OpenStudy (anonymous):

Find the y-intercept of this linear equation. y - 4x = 12

OpenStudy (goten77):

again.. just plug in 0 for x y - 0 = 12 y = 12 is the intercept (0,12)

OpenStudy (ashley1nonly):

y-4x=12 ad the 4x to both sides now you have y=4x+12 which is in y=mx+b so (0,12)